X squared = x+42
Confused and not able to come up with an answer that works. I would appreciate your help with this one.

Answer by mananth(16946)

x^2=x+42
x^2-x-42=0
x^2-7x+6x-42=0
x(x-7)+6(x-7)=0
(x+6)(x-7)=0
x=-6,OR x=7